# Can AR Embedded Visualizations Foster Appropriate Reliance on AI in Spatial Decision-Making? A Comparative Study of AR X-Ray vs. 2D Minimap

### v3

@@ 5.2. Qualitative Results
We conducted semi-structured post-study interviews to capture participants’ strategies, rationales, and motivations behind their decision-making across all conditions. Participants are referred to as P1–P32. We analyzed the interviews using reflexive thematic analysis (braun2019reflecting), identifying themes both inductively (bottom-up) and deductively (top-down). Below, we summarize the key themes that emerged from the analysis.
@@ 5.2.1. Diverse Decision Strategies Across Visualizations
[…]
@@ 5.2.2. Increased Trust in AI Suggestions with X-Ray
  • Perceptual and Navigation Challenges. Participants found distance estimation (P2), queue length assessment (P4, P5), and locating navigation paths (P12) more effortful in AR. Thus, AI suggestions offered a useful cognitive shortcut, aligning with findings from prior research on AI-aided decision-making (2021BansalDoesTheWholeExceedItsParts).
@@ 6.1. Perceptual Challenges of AR X-Ray Can Lead to Over-Reliance on AI
Our findings suggest that perceptual challenges of the AR X-ray, such as occlusion, large canvas space, and spatial estimation difficulty, can increase users’ reliance on AI suggestions, even when those suggestions are suboptimal. This over-reliance does not appear to stem from blind trust in AI, but rather from the inherent complexity of perceiving and evaluating visual information in large-scale physical environments. While prior research has discussed AR’s perceptual limitations (Drascic1996PerceptualII), our study directly connects these challenges to AI reliance patterns in time-sensitive spatial decision-making. This highlights the need for better embedded visualization designs that support spatial reasoning, especially in large-scale environments.
@@
Future Studies. In our study, situational data like queue length was visualized using human-shaped glyphs. A natural possibility is to compare visual encoding, for example, glyphs vs. numeric labels, to evaluate how data representation like legibility, uncertainty and saliency influences AI reliance.
@@ 6.2. Cognitive Biases Triggered by AR X-Ray Can Induce Inappropriate AI Reliance
Beyond perceptual difficulty, our qualitative findings point to cognitive biases that are unique to the AR X-ray. Participants reported visual proximity illusions and increased trust in AI suggestions due to their spatial embodiment in the environment. These effects align with prior research on spatial cognition biases (raghubir1996crow) and trust in embodied agents (Kim2018DigitalAssistantNeedABody).
Although cognitive biases have been studied in traditional visualization contexts (dimara2018task; wall2021left), their implications in embodied, spatial, and AI-augmented AR systems remain underexplored. As Padilla et al. (padilla2018decision) suggested, when systems are designed with such biases in mind, users can benefit from Type 1 (intuitive) thinking—particularly under time pressure. Our findings highlight a fertile area for interdisciplinary research at the intersection of spatial cognition, embedded visualizations, and human-AI decision-making.
@@ 6.3. Aligning the Strengths of AR X-Ray with Spatial Tasks
Design Implications. AR embedded visualizations may be most effective when tightly coupled with tasks involving physical action, rather than static decision-making. Designers should consider the viewer’s movement and real-time orientation when designing the embedded visualizations. Emerging work on “Visualization in Motion” (DBLP:journals/IEEE_TVCG/YaoBVI22) and “First-Person View Visualization” (isenberg:hal-04782928) provides valuable guidance in this direction. In addition, AR systems should support smooth transitions between embedded visualizations and traditional 2D visualizations in the headset depending on task context and user needs.
@@ 6.4. AR X-Ray May Lower Cognitive Engagement Compared with 2D Visualizations
In our study, there was no significant difference in response time between the Minimap+NoAI and X-ray+NoAI conditions once initial search time was excluded. However, participants in the X-ray+AI condition made decisions more quickly yet achieved lower overall accuracy than those in the Minimap+AI condition. This pattern aligns with reliance on heuristic judgments, often described as Type 1 thinking (Chaiklin2012ThinkingFA; JSBTDualProcessTheory2013). Together with the high self-reported confidence, these results suggest a lack of deeper cognitive engagement: once AI assistance was available, participants quickly felt confident enough in their intuitive judgments, rather than engaging in more analytical (Type 2) thinking. These findings imply that AR X-Ray, by offering a highly immersive or embodied experience, may inadvertently reduce users’ motivation to scrutinize information.
Design Implications. It remains unclear whether this lower cognitive engagement is specific to our task design, a broader characteristic of AR embedded visualizations, or a more general cognitive shift from 2D to AR egocentric perspective. Design strategies that prompt more deliberation, such as “cognitive forcing functions” (2021ToTrustOrToThink), could be incorporated to encourage deeper analytic thinking in AR environments.
@@ 6.5. Study Limitations
Our study offers an initial examination of AI-aided spatial decision-making with AR embedded visualizations. Although guided by prior research on AR visual systems for spatial tasks (XUImroveIndoorWayfinding; XUARteambasedsearch), we acknowledge that our system design does not represent a fully optimized, advanced, or universal solution.
To ensure experimental control, we adopted fixed parameters, such as a discrete set of spatial arrangements, an AI accuracy of 75% (see Section 4.7), and a 20-second decision window, which may not capture the full complexity of high-stakes, real-world environments or more advanced AI systems. Consequently, our findings are most transferable to tasks resembling the indoor, time-sensitive context described in Section 3.1. Moreover, our participant pool primarily comprised well-educated individuals, which may not reflect the behavior of specialized user groups, for example, emergency responders or security professionals. As is common in spatial cognition research (Tan2024InvisibleME), our sample size was modest; although sufficient for controlled experimentation, a larger and more diverse sample would strengthen external validity.

### v4

@@ 5.2. Qualitative Results
We conducted semi-structured post-study interviews to capture participants’ strategies, rationales, and motivations behind their decision-making across all conditions. Participants are referred to as P1–P32. We analyzed the interviews using reflexive thematic analysis (Braun and Clarke, 2019), identifying themes both inductively (bottom-up) and deductively (top-down). Below, we summarize the key themes that emerged from the analysis.
@@ 5.2.1. Diverse Decision Strategies Across Visualizations
Example of occlusion in the AR X-ray visualization, where overlapping avatars make it difficult to estimate the number of people in the queue, leading to potential misjudgment of wait time.
@@ 5.2.2. Increased Trust in AI Suggestions with X-Ray
  • Perceptual and Navigation Challenges. Participants found distance estimation (P2), queue length assessment (P4, P5), and locating navigation paths (P12) more effortful in AR. Thus, AI suggestions offered a useful cognitive shortcut, aligning with findings from prior research on AI-aided decision-making (Bansal et al., 2021).
@@ 6.1. Perceptual Challenges of AR X-Ray Can Lead to Over-Reliance on AI
Our findings suggest that perceptual challenges of the AR X-ray, such as occlusion, large canvas space, and spatial estimation difficulty, can increase users’ reliance on AI suggestions, even when those suggestions are suboptimal. This over-reliance does not appear to stem from blind trust in AI, but rather from the inherent complexity of perceiving and evaluating visual information in large-scale physical environments. While prior research has discussed AR’s perceptual limitations (Drascic and Milgram, 1996), our study directly connects these challenges to AI reliance patterns in time-sensitive spatial decision-making. This highlights the need for better embedded visualization designs that support spatial reasoning, especially in large-scale environments.
@@
Future Studies. In our study, situational data like queue length was visualized using human-shaped glyphs. A natural possibility is to compare visual encoding (e.g., glyphs vs. numeric labels) to evaluate how data representation like legibility, uncertainty and saliency influences AI reliance.
@@ 6.2. Cognitive Biases Triggered by AR X-Ray Can Induce Inappropriate AI Reliance
Beyond perceptual difficulty, our qualitative findings point to cognitive biases that are unique to the AR X-ray. Participants reported visual proximity illusions and increased trust in AI suggestions due to their spatial embodiment in the environment. These effects align with prior research on spatial cognition biases (Raghubir and Krishna, 1996) and trust in embodied agents (Kim et al., 2018).
Although cognitive biases have been studied in traditional visualization contexts (Dimara et al., 2018; Wall et al., 2021), their implications in embodied, spatial, and AI-augmented AR systems remain underexplored. As Padilla et al. (Padilla et al., 2018a) suggested, when systems are designed with such biases in mind, users can benefit from Type 1 (intuitive) thinking—particularly under time pressure. Our findings highlight a fertile area for interdisciplinary research at the intersection of spatial cognition, embedded visualizations, and human-AI decision-making.
@@ 6.3. Aligning the Strengths of AR X-Ray with Spatial Tasks
Design Implications. AR embedded visualizations may be most effective when tightly coupled with tasks involving physical action, rather than static decision-making. Designers should consider the viewer’s movement and real-time orientation when designing the embedded visualizations. Emerging work on “Visualization in Motion” (Yao et al., 2022) and “First-Person View Visualization” (Isenberg et al., 2024) provides valuable guidance in this direction. In addition, AR systems should support smooth transitions between embedded visualizations and traditional 2D visualizations in the headset depending on task context and user needs.
@@ 6.4. AR X-Ray May Lower Cognitive Engagement Compared with 2D Visualizations
In our study, there was no significant difference in response time between the Minimap+NoAI and X-ray+NoAI conditions once initial search time was excluded. However, participants in the X-ray+AI condition made decisions more quickly yet achieved lower overall accuracy than those in the Minimap+AI condition. This pattern aligns with reliance on heuristic judgments, often described as Type 1 thinking (Chaiklin, 2012; Evans and Stanovich, 2013). Together with the high self-reported confidence, these results suggest a lack of deeper cognitive engagement: once AI assistance was available, participants quickly felt confident enough in their intuitive judgments, rather than engaging in more analytical (Type 2) thinking. These findings imply that AR X-Ray, by offering a highly immersive or embodied experience, may inadvertently reduce users’ motivation to scrutinize information.
Design Implications. It remains unclear whether this lower cognitive engagement is specific to our task design, a broader characteristic of AR embedded visualizations, or a more general cognitive shift from 2D to AR egocentric perspective. Design strategies that prompt more deliberation, such as “cognitive forcing functions” (Buçinca et al., 2021), could be incorporated to encourage deeper analytic thinking in AR environments.
@@ 6.5. Study Limitations
Our study offers an initial examination of AI-aided spatial decision-making with AR embedded visualizations. Although guided by prior research on AR visual systems for spatial tasks (Xu et al., 2024b, a), we acknowledge that our system does not represent a fully optimized, advanced, or universal solution.
To ensure experimental control, we adopted fixed parameters, such as a discrete set of spatial arrangements, an AI accuracy of 75% (see Section 4.7), and a 20-second decision window, which may not capture the full complexity of high-stakes, real-world environments or more advanced AI systems. Consequently, our findings are most transferable to tasks resembling the indoor, time-sensitive context described in Section 3.1. Moreover, our participant pool primarily comprised well-educated individuals, which may not reflect the behavior of specialized user groups, for example, emergency responders or security professionals. As is common in spatial cognition research (Tan et al., 2024), our sample size was modest; although sufficient for controlled experimentation, a larger and more diverse sample would strengthen external validity.